OpenStudy (anonymous):

find all solutions to tan x=1

OpenStudy (anonymous):

\[\frac{ sinx }{ cosx }=1\]

OpenStudy (anonymous):

in terms of n

OpenStudy (anonymous):

what do you have in mind

OpenStudy (anonymous):

i can go from there

OpenStudy (anonymous):

\[\tan x=1\] \[ x=\frac{\pi}{4}, \frac{5\pi}{4}\]